OpenStudy (anonymous):

Solve. 3(x – 2) + 5x = 18 (Points : 4) all real numbers 1 2 3

OpenStudy (anonymous):

is it 3?

OpenStudy (anonymous):

Distribute \(3\) inside the brackets.. :)

OpenStudy (igreen):

3(x - 2) + 5x = 18 Distribute: 3x - 6 + 5x = 18 Simplify: 8x - 6 = 18 Add 6 to both sides: 8x = 24 Divide 8 to both sides, can you do that? @marie123

OpenStudy (anonymous):

Yep it is \(3\), Good.. :)

OpenStudy (igreen):

Yep, you got it. @marie123